OpenStudy (anonymous):

find an equation of the line containing the given pair of points (-2,-2) and (-4,-9) y=

OpenStudy (anonymous):

so you have pt 1 (-2,-2) (x1,y1) and pt 2 (-4,-9), (x2,y2). you need to find the slope (m) of the line by m= (y2 - y1) / (x2 - x1). So then m = (-9 - (-2)) / (-4 - (-2)). m = -7/-2, m = 7/2. Then you plug that into the line equation y = mx+b. So now y = (7/2)x +b. If you take the pt (-2,-2) and plug those x, and y values into the same equation, you can solve for b, which is the y-intercept of that line. -2 = (7/2)(-2) + b. 5 = b. so the line with those points is satisfied by y = (7/2)x + 5. All other questions like this can be approached the same way.